# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

### v3

@@ 4 Conclusions
We show that interpolating neural networks (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods or machine learning models. The keys to INN’s success in those problems are 1) it interpolates virtual points with well-established interpolation theories (not interpolating training data with a black-box network) and 2) it leverages the idea of separation of variables and tensor decomposition to resolve the curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INN, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Multi-resolution aspect (i.e., a varying number of segments across INN networks, $M$) of INN needs to be investigated. See Supplementary Information for further discussions.
  • Convergence of INN for training can be proved mathematically.
  • INN training on ensembles of training data might be a way to enhance the generality of the model for image problems or noisy data.
  • Since INN could be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors together into one single model, so that we can solve large scale problems involving both complex physics and scarce data. For example, the calibration task of LPBF heat source.
We expect INN can open a next-generation Engineering Software 2.0 by theoretically integrating training, solving, and calibration and tackling extremely challenging problems in the science and engineering domain.

### v4

@@ 4 Conclusions
This article demonstrates that the interpolating neural network (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models. The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories (not interpolating training data with a black-box network that often causes overfitting issues) and 2) it leverages the tensor decomposition (TD) to resolve curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INNs, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Depending on the problem, INN’s superior training efficiency may cause overfitting issues. The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the model for image problems or noisy data.
  • Multi-resolution aspect (i.e., a varying number of segments across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
  • Similar to the INN solvers, the convergence of INN trainers can be studied mathematically.
  • Since INNs can be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors into one single model to solve large-scale problems involving both complex physics and scarce data.
  • The superior performance of the INN solver may facilitate multiscale analysis in a vast parametric space. One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [53, 54] and open a new direction towards parameterized multiscale analysis.
  • The current INN code has been optimized to run on a single GPU. Since the INN forward pass over multiple modes can be parallelized, multi-GPU programming will further speed up the code.
  • Although Tucker decomposition has been introduced as a special case 1 in Section 2, the numerical experiments conducted in this article mostly focus on the CP decomposition. Theoretically, Tucker decomposition has a larger approximation space than CP decomposition. Further studies on the numerical aspects of the Tucker decomposition are needed.
We expect INNs will open a next-generation Engineering Software 2.0 that revolutionizes almost all fields of computational science and engineering.
Acknowledgments
We express our sincere gratitude to Dr. Gino Domel, Mr. Joseph P. Leonor, Mr. Stefan Knapik, and Mr. Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work. We also acknowledge the following grants: DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-20-2-0292 and W911NF-21-2-02199 to J. Cao and the startup funds provided by Virginia Polytechnic Institute and State University to S. Saha. C. Park acknowledges the internship opportunity with the Freudenberg Group, within the team managed by Dr. Vishwanath Hegadekatte.